You pull a box using a rope, which is directed 30 degrees above the horizontal. You apply a constant force of 30 N, and you move the box 5 m along the ground. What work did you do in Joules?

1
Expert's answer
2022-04-21T17:44:57-0400

By definition, the work is the scalar product of the force and displacement:


"W = \mathbf{F}\cdot \mathbf{d}"

Since the box is displaced along the horizontal direction, and the force makes "\theta=30\degree" with horizontal, the formula for the scalar product becomes:


"W = Fd\cos\theta"

where "F = 30N, d = 5m". Thus, obtain:


"W = 30N\cdot 5m\cdot \cos30\degree \approx 130J"

Answer. 130 J.